OpenStudy (anonymous):

At a physics convention, 10 companies set up equal sized square booths in a row along one wall of the convention center. The booths are adjacent to each other and a 4-ft wide walkway surrounds the block of booths on three sides. The total area of the booths and walkway is 2300 ft^2. What is the side length of each booth?

OpenStudy (anonymous):

set up the problem: \[(x+4)(10x+8)=2300\] \[10x^2+48x+32=2300\] \[10x^2+48x-2268=0\] \[(-48\pm \sqrt{48^2-4*10*(-2268)})/(2*10)\] x~12.85